OpenStudy (anonymous):

multiply (2x+5)(2x+5)

OpenStudy (anonymous):

nikkiem1 please give us an example of hint of what variable x is.

OpenStudy (anonymous):

4x^2 +14x+25

OpenStudy (anonymous):

there is no hint of the variable that is where i was lost too

OpenStudy (anonymous):

Qed so the answer then would be 56x+25 right?

OpenStudy (anonymous):

i mean 30x+25???

OpenStudy (anonymous):

(a+b)^2=a^2+2ab+b^2 4x^2 + 20x + 25

OpenStudy (anonymous):

the product is 4x^2+20x+25

OpenStudy (anonymous):

so then what would be the answer i mean what do i do times 4x to the 2 power by 20 then add 25

OpenStudy (anonymous):

use FOIL method

OpenStudy (anonymous):

what is that

OpenStudy (anonymous):

multiply the expressions, first term of both expressions, then outer terms, then inner terms, and last terms.

OpenStudy (anonymous):

Look there http://www.freemathhelp.com/using-foil.html

OpenStudy (anonymous):

so it looks like this (2x)(2x)+(5)(2x)+(5)(2x)+(5)(5)

OpenStudy (anonymous):

i times 4x to the 2nd power by 20 i got 320 am i right so far

OpenStudy (anonymous):

2x*2x=4x^2

OpenStudy (anonymous):

so your the answer then would be 49 right?

OpenStudy (anonymous):

you don't need to expand the answer, you just need to multiply the expressions. do you get it???

OpenStudy (anonymous):

im so lost

OpenStudy (anonymous):

so then it would be 16

OpenStudy (anonymous):

what grade are you in??

OpenStudy (anonymous):

Your final answer is 4x^2 + 20x + 25 Look at both link. It will help you to understand what you are doing. http://www.coolmath.com/prealgebra/15-intro-to-polynomials/06-polynomials-multiplying-foil-01-108.htm

OpenStudy (anonymous):

that seem to help a little thank you

OpenStudy (anonymous):

You are welcome. It really will help you to practice FOIL on your own. After spending 10 minutes on that you will be able to answer most of you today's questions by yourself.